# Active Measurement: Efficient Estimation at Scale

### v1

@@ B.1 The plug-in mean in variance estimation
Given the same assumptions as Prop. 3, with at least probability $1-p$, there exists a constant $H$ dependent on $p$ such that $|\widetilde{\operatorname{\mathrm{Var}}}_{\tau,t}-\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau,t}|\leq H(N-\tau+1)\sqrt{\frac{N-t+1}{t-1}}$.
@@
By Chebyshev’s inequality, with at least probability $1-p$,
@@ B.2 The streaming algorithm of variance estimation
Alg. 2 has a summation for each $1\leq\tau\leq t$, amounting to a complexity of $\mathcal{O}(t^{2})$. In practice, we would reuse part of previous computation at step $t-1$ to accelerate the computation at step $t$. Observe that
@@ C.1 Estimating birds in radar data
Fig. A2 (left) illustrates weather radar scans capturing bird roosts. Fig. A2 (right) visualizes the reflectivity data at 0.50.5$\circ$0.5 ∘ elevation of a radar scan that happened at the KCLE weather radar station on August 19 2015 at 10:31:22 UTC, which is a randomly sampled radar scan that captures bird roosts. The circular and semicircular patterns are massive amounts of birds departing from their overnight roosting locations.
@@
  • Per-bounding-box counts. We summarize, for each detection, radar sweeps taken across multiple elevations. In order to prevent double counting of birds in regions sampled twice by two consecutive beams, we bin the sweep elevations into 11$\circ$1 ∘ bins. We then take the average of sweeps that fall in the same bin. Lastly, we sum counts across all bins.
@@ C.2 Additional Finetuning And Sampling Details
In practice, the measurements from the detectors could be as low as 00, making the labeling of some units impossible and introducing huge variances to the estimation. In our image counting tasks, we set $g(s)\leftarrow\max(g(s),1)$ to make sure that each tile can be sampled. In our radar counting tasks, we set $g(s)\to g(s)+1000$ to cover every day in every station.
@@ D.5 Confidence intervals for all experiments
D.6 More comparisons with other baselines
We also compare active measurement with two other baselines: active testing and $\hat{H}_{t}$ (motivated by PPI), on the image counting tasks. The results are in Fig. A8. In general, we find active measurement has the lowest error.
Figure A8: Fractional error compared with other baselines, for the two image counting tasks, averaged over 10,000 simulations.

### v2

@@ B.1 The plug-in mean in variance estimation
Given the same assumptions as Prop. 3, with at least probability 1p1-p, there exists a constant $H$ dependent on $p$ such that $|\widetilde{\operatorname{\mathrm{Var}}}_{\tau,t}-\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau,t}|\leq H(N-\tau+1)\sqrt{\frac{N-t+1}{t-1}}$.
@@
By Chebyshev’s inequality, with at least probability 1p1-p,
@@ B.2 The streaming algorithm of variance estimation
Alg. 2 has a summation for each $1\leq\tau\leq t$, amounting to a complexity of $\mathcal{O}(t^{2})$. In practice, we would reuse part of previous computation at step t1t-1 to accelerate the computation at step $t$. Observe that
@@ C.1 Estimating birds in radar data
Fig. A2 (left) illustrates weather radar scans capturing bird roosts. Fig. A2 (right) visualizes the reflectivity data at $0.5\circ$ elevation of a radar scan that happened at the KCLE weather radar station on August 19 2015 at 10:31:22 UTC, which is a randomly sampled radar scan that captures bird roosts. The circular and semicircular patterns are massive amounts of birds departing from their overnight roosting locations.
@@
  • Per-bounding-box counts. We summarize, for each detection, radar sweeps taken across multiple elevations. In order to prevent double counting of birds in regions sampled twice by two consecutive beams, we bin the sweep elevations into $1\circ$ bins. We then take the average of sweeps that fall in the same bin. Lastly, we sum counts across all bins.
@@ C.2 Additional Finetuning And Sampling Details
In practice, the measurements from the detectors could be as low as 0, making the labeling of some units impossible and introducing huge variances to the estimation. In our image counting tasks, we set $g(s)\leftarrow\max(g(s),1)$ to make sure that each tile can be sampled. In our radar counting tasks, we set $g(s)\to g(s)+1000$ to cover every day in every station.